OpenStudy (anonymous):

Sum upto infinity of divergent series Usually i had thought when we have divergent series the sum upto infinity is infinity.. but 1+2+3+4... = -1/12 :P.. now i looked this up, and of course there are proofs available, but logic disproves it, like for eg i can use simple mathematics and say 1+2+3.. > 1+2.. so all those things say the result is wrong.. also i can do 1+2+3+4 ... = -1/12 1+ (1+1)+(1+2)+(1+3)= -1/12 (1+1+1+1....) +(1+2+3+ ... ) = -1/12 (1+1+1+1..)+ (-1/12)= -1/12 hence (1+1+1+1.....) = 0 so i don't want u guys to prove it :p..

OpenStudy (anonymous):

can someone help me understand why our DISPROOFS using simple maths is invalid? @Vincent-Lyon.Fr @AllTehMaffs @phi @cwrw238 @ajprincess @Compassionate

OpenStudy (atlas):

your assumption 1+2+3+4+....=-1/12 is invalid

OpenStudy (anonymous):

its not my assumption @atlas its proved to be true.. believe me.. i don't believe it.. but i did research and everywhere its true..

OpenStudy (anonymous):

here is one source numberphile :D http://kottke.org/14/01/the-sum-of-all-positive-integers

OpenStudy (atlas):

it is invalid

OpenStudy (anonymous):

:-/ could you elaborate? or go through the site.. before just bluntly saying its invalid? :P.. yes i know it looks invalid..

OpenStudy (atlas):

let me check put

OpenStudy (anonymous):

thank you :D

OpenStudy (atlas):

the site does not open up here....can you give me any other link

OpenStudy (anonymous):

ok here is a wiki link http://en.wikipedia.org/wiki/1_%2B_2_%2B_3_%2B_4_%2B_%E2%8B%AF

OpenStudy (anonymous):

This formula actually exists in a book, like the guy in the video said. However, we all know this sum is incorrect because it appears in a physics book XD

OpenStudy (anonymous):

The person in the video is adding two divergent series terms by terms. You cannot do that. In fact you can arrange any convergent alternating series to converge to any number.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Alternating_series

OpenStudy (anonymous):

Say how they prove in the above link that 1=0

OpenStudy (atlas):

ok i saw the proof and it seemed mathematically correct but the proof was obtained by rearranging the terms of a divergent series which did not feel right....:

OpenStudy (atlas):

It is like saying: S = 1+2+3+4+.... S= 1+2+3+...... Subtracting both 0 = 1+1+1+1...... and so 0 = infinity

OpenStudy (turingtest):

Indeed, this answer from Quora agrees with @eliassaab http://www.quora.com/Mathematics/Theoretically-speaking-how-can-the-sum-of-all-positive-integers-be-1-12 They explain that zeta(-1)/12 but that's a result of something I don't seem to understand.

OpenStudy (anonymous):

ok guys thanks for taking interest.. i have realized that the word "summation" itself is at stake here, so this summation is not like adding 1 apple + 2 apple + 3 apple +.... upto infinity .. cause if we did.. we know we can't get -1/12 apples :D.. So i guess this is something that will be applied in string theory and quantum mechanics only.. ll do further research to see it.. thanks again guys :D

OpenStudy (anonymous):

@eliassaab that i think was just a simple proof.. there are more elaborate proofs.. but my original question was not to understand the proof, but to understand the meaning :P

OpenStudy (anonymous):

@Mashy, in logic if statement A is false then for any statement B, the statement \[ A \implies B \] is true. If you use a false method, you can prove anything you want.

OpenStudy (anonymous):

no no.. they just used that method in laymans terms.. it has been proved using some complicated math.. look up Ramanujan series.. the great indian scientist has ALSO proved it.. its called as Ramanujan summation.. so don't bother about the VALIDITY.. or its proof.. its right.. its been accepted officially.. i just wanted to understand it!

OpenStudy (anonymous):

I am hoping that you can understand it. Good luck.

OpenStudy (anonymous):

thank you :)

OpenStudy (atlas):

@Mashy but I understood your question and it was a good question to ask. However, I am not so deep into mathematics but I hope you get the right person to answer the question. Besides I think this link might be useful http://math.arizona.edu/~cais/Papers/Expos/div.pdf and you may like to follow it up :)